# DiffusionGAN3D: Boosting Text-guided 3D Generation and Domain Adaption by Combining 3D GANs and Diffusion Priors

### 29 Dec 2023 v2

@@ Abstract
Text-guided domain adaption and generation of 3D-aware portraits find many applications in various fields. However, due to the lack of training data and the challenges in handling the high variety of geometry and appearance, the existing methods for these tasks suffer from issues like inflexibility, instability, and low fidelity. In this paper, we propose a novel framework DiffusionGAN3D, which boosts text-guided 3D domain adaption and generation by combining 3D GANs and diffusion priors. Specifically, we integrate the pre-trained 3D generative models (e.g., EG3D) and text-to-image diffusion models. The former provides a strong foundation for stable and high-quality avatar generation from text. And the diffusion models in turn offer powerful priors and guide the 3D generator finetuning with informative direction to achieve flexible and efficient text-guided domain adaption. To enhance the diversity in domain adaption and the generation capability in text-to-avatar, we introduce the relative distance loss and case-specific learnable triplane respectively. Besides, we design a progressive texture refinement module to improve the texture quality for both tasks above. Extensive experiments demonstrate that the proposed framework achieves excellent results in both domain adaption and text-to-avatar tasks, outperforming existing methods in terms of generation quality and efficiency. The project homepage is at https://younglbw.github.io/DiffusionGAN3D-homepage/.
@@ 1 Introduction
Some works [56, 2, 51, 53, 25, 26, 47] perform transfer learning on a pre-trained 3D GAN generator to achieve 3D stylization, which relies on a large number of stylized images and strictly aligned camera poses for training. [2, 47] leverage existing 2D-GAN trained on a specific domain to synthesize training data and implement finetuning with adversarial loss. In contrast, [51, 25, 26] utilize text-to-image diffusion models to generate training datasets in the target domain. This enables more flexible style transferring but also brings problems like pose bias, tedious data processing, and heavy computation costs. Unlike these adversarial finetuning based methods, StyleGAN-Fusion [48] adopts SDS [37] loss as guidance of text-guided adaption of 2D and 3D generators, which gives a simple yet effective way to fulfill domain adaption. However, it also suffers from limited diversity and suboptimal text-image correspondence.
@@
Based on the observations above, we propose a novel two-stage framework DiffusionGAN3D to boost the performance of 3D domain adaption and text-to-avatar tasks by combining 3D generative models and diffusion priors, as shown in Fig. 2. For the text-guided 3D Domain Adaption task, We first leverage diffusion models and adopt SDS loss to finetune a pre-trained EG3D-based model [7, 4, 9] with random noise input and camera views. The relative distance loss is introduced to deal with the loss of diversity caused by the SDS technique. Additionally, we design a diffusion-guided reconstruction loss to adapt the framework to local editing scenarios. Then, we extend the framework to Text-to-Avatar task by finetuning 3D GANs with a fixed latent code that is obtained guided by CLIP [38] model. During optimization, a case-specific learnable triplane is introduced to strengthen the generation capability of the network. To sum up, in our framework, the diffusion models offer powerful text-image priors, which guide the domain adaption of the 3D generator with informative direction in a flexible and efficient way. In turn, 3D GANs provide a strong foundation for text-to-avatar, enabling stable and high-quality avatar generation. Last but not least, taking advantage of the powerful 2D synthesis capability of diffusion models, we propose a Progressive Texture Refinement module as the second stage for these two tasks above, which significantly enhances the texture quality. Extensive experiments demonstrate that our method exhibits excellent performance in terms of generation quality and stability on 3D domain adaption and text-to-avtar tasks, as shown in Fig. 1.
@@
(A) We achieve text-guided 3D domain adaption in high quality and diversity by combining 3D GANs and diffusion priors with the assistance of the relative distance loss.
@@ 2 Related Work
Domain Adaption of 3D GANs. The advancements in 3D generative models [7, 35, 4, 9, 6, 11, 14, 13, 29, 45] have enabled geometry-aware and pose-controlled image generation. Especially, EG3D [7] utilizes triplane as 3D representation and integrates StyleGAN2 [24] generator with neural rendering [33] to achieve high-quality 3D shapes and view-consistency image synthesis, which facilitates the downstream applications such as 3D stylization, GAN inversion [28]. Several works [22, 56, 2, 53] achieve 3D domain adaption by utilizing stylized 2D generator to synthesize training images or distilling knowledge from it. In contrast, [25, 26, 51] leverage the powerful diffusion models to generate training datasets in the target domain and accomplish text-guided 3D domain adaption with great performance. Though achieving impressive results, these adversarial learning based methods above suffer from issues such as pose bias, tedious data processing, and heavy computation cost. Recently, non-adversarial finetuining methods [3, 12, 48] also exhibit great promise in text-guided domain adaption. Especially, StyleGAN-Fusion [48] adopts SDS loss as guidance for the adaption of 2D generators and 3D generators. It achieves efficient and flexible text-guided domain adaption but also faces the problems of limited diversity and suboptimal text-image correspondence.
@@ 3 Methods
In this section, we present DiffusionGAN3D, which boosts the performance of 3D domain adaption and text-to-avatar by combining and taking advantage of 3D GANs and diffusion priors. Fig. 2 illustrates the overview of our framework. After introducing some preliminaries (Sec. 3.1), we first elaborate our designs in diffusion-guided 3D domain adaption (Sec. 3.2) , where we propose a relative distance loss to resolve the problem of diversity loss caused by SDS. Then we extend this architecture and introduce a case-specific learnable triplane to fulfill 3D-GAN based text-to-avatar (Sec. 3.3). Finally, we design a novel progressive texture refinement stage (Sec. 3.4) to improve the detail and authenticity of the texture generated above.
@@ 3.1 Preliminaries
3.2 Diffusion-Guided 3D Domain Adaption
Due to the difficulties in obtaining high-quality pose-aware data and model training, adversarial learning methods for 3D domain adaption mostly suffer from the issues of tedious data processing and mode collapse. To address that, we leverage diffusion models and adopt the SDS loss to implement transfer learning on an EG3D-based 3D GAN to achieve efficient 3D domain adaption, as shown in Fig. 2.
Given a style code $\bm{w}$ generated from noise $\bm{z}\sim N(0,1)$ through the fixed mapping network, we can obtain the triplane $\bm{T}$ and the image $\bm{x}$ rendered in a view controlled by the input camera parameters $\bm{c}$ using the triplane generator and decoder in sequence. Then SDS loss (Sec. 3.1) is applied on $\bm{x}$ to finetune the network. Different from DreamFusion which optimizes a NeRF network to implement single object generation, we shift the 3D generator with random noise and camera pose to achieve domain adaption guided by text $\bm{y}$. During optimization, all parameters of the framework are frozen except the triplane generator. We find that the gradient provided by SDS loss is unstable and can be harmful to some other well-trained modules such as the super-resolution module. Besides, freezing the mapping network ensures that the latent code $\bm{w}$ lies in the same domain during training, which is a crucial feature that can be utilized in the diversity preserving of the 3D generator.
Relative Distance Loss. The SDS loss provides diffusion priors and achieves text-guided domain adaption of 3D GAN in an efficient way. However, it also brings the problem of diversity loss as illustrated in [48]. To deal with that, [48] proposes the directional regularizer to regularize the generator optimization process, which improves the diversity to a certain extent. However, it also limits the domain shifting, facing a trade-off between diversity and the degree of style transfer. To address this, we propose a relative distance loss. As shown in Fig. 3, considering two style codes $\bm{w_{i}}$ and $\bm{w_{j}}$ which are mapping from two different noise $\bm{z_{i}}$ and $\bm{z_{j}}$, we project them into the original triplane domain ($\bm{T^{\prime}_{i}}$, $\bm{T^{\prime}_{j}}$) and the finetuned one ($\bm{T_{i}}$, $\bm{T_{j}}$) using a frozen triplane generator $G_{frozen}$ and the finetuned triplane generator $G_{train}$, respectively. Note that, since the mapping network is frozen during training in our framework, $\bm{T_{i}}$ and $\bm{T^{\prime}_{i}}$ (same for $\bm{T_{j}}$ and $\bm{T^{\prime}_{j}}$) share the same latent code and ought to be close in context. Thus, we model the relative distance of these two samples in triplane space and formulate the relative distance loss $L_{dis}$ as:
@@ 3.2 Diffusion-Guided 3D Domain Adaption
Diffusion-guided Reconstruction Loss. Despite the combination of SDS loss and the proposed relative distance loss is adequate for most domain adaption tasks, it still fails to handle the local editing scenarios. A naive solution is to perform reconstruction loss between the rendered image and the one from the frozen network. However, it will also inhibit translation of the target region. Accordingly, we propose a diffusion-guided reconstruction loss especially for local editing, which aims to preserve non-target regions while performing 3D editing on the target region. We found that the gradient of SDS loss has a certain correlation with the target area, especially when the noise level $t$ is large, as shown in Fig. 4. To this end, we design a diffusion-guided reconstruction loss $L_{diff}$ that can be presented as:
@@
To sum up, we can form the loss functions for normal domain adaption and local editing scenario as $L_{adaption}=L_{sds}+\lambda_{1}L_{dis}$ and $L_{editing}=L_{sds}+\lambda_{2}L_{diff}$, respectively, where $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients.
@@ 4.1 Implementation Details
Our framework is built on an EG3D-based model in the first stage. Specifically, we implement 3D domain adaption on PanoHead, EG3D-FFHQ, and EG3D-AFHQ for head, face, and cat, respectively. For text-to-avatar tasks, PanoHead and AG3D are adopted as the base generators for head and body generation. We employ StableDiffusion v2.1 as our pre-trained text-to-image model. In the texture refinement stage, StableDiffusion v1.5 coupled with ControlNets are utilized to implement image-to-image and inpainting. More details about the parameters and training setting are specified in supplementary materials.
@@ 4.2 Qualitative Comparison
Figure 6: The qualitative comparisons on 3D domain adaption (applied on EG3D-FFHQ [7]).
For 3D Domain adaption, we evaluate our model with two powerful baselines: StyleGAN-NADA* [12] and StyleGAN-Fusion [48] for text-guided domain adaption of 3D GANs, where * indicates the extension of the method to 3D models. For a fair comparison, we use the same prompts as guidance for all the methods. Besides, the visualization results of different methods are sampled from the same random noise. As shown in Fig. 6, the naive extension of StyleGAN-NADA* for EG3D exhibits poor results in terms of diversity and image quality. StyleGAN-Fusion achieves decent 3D domain adaption and exhibits a certain diversity. However, the proposed regularizer of StyleGAN-Fusion also hinders itself from large-gap domain transfer, resulting in a trade-off between the degree of stylization and diversity. As Fig. 6 shows that the generated faces of StyleGAN-Fusion lack diversity and details, and the hair and clothes suffer from inadequate stylization. In contrast, our method exhibits superior performance in diversity, image quality, and text-image correspondence.
Figure 7: Visual comparisons on text-to-avatar task. The first two rows are the results of ’head’ and the rest are the results of ’body’.
@@ 4.3 Quantitative Comparison
We quantitatively evaluate the above baselines and our method on 3D domain adaption through FID [15] comparison and user study. Specifically, all methods are employed to conduct domain adaption on EG3D-face and EG3D-cat with both four text prompts, respectively. For each text prompt, we utilize the text-to-image diffusion model to generate 2000 images with different random seeds as the ground truth for FID calculation. In the user study, 12 volunteers were invited to rate each finetuned model from 1 to 5 based on three dimensions: text-image correspondence, image quality, and diversity. As shown in Table  1, the proposed methods achieve lower FID scores than other baselines, which indicates superior image fidelity. Meanwhile, the user study demonstrates that our method outperforms the other two methods, especially in terms of image quality and diversity.
@@
Table 1: Quantitative comparison on 3D domain adaption task.
@@ 4.4 Ablation Study
On relative distance loss. As shown in Fig. 9, if adopting SDS loss alone for domain adaption, the generator will tend to collapse to a fixed output pattern, losing its original diversity. In contrast, the proposed relative distance loss effectively preserves the diversity of the generator without sacrificing the stylization degree.
@@
Figure 10: Ablation study of the diffusion guided reconstruction loss. The ToRGB module in EG3D is trained together with $G_{train}$. The input text is ”a close-up of a woman with green hair”.
@@ 5 Conclusion
In this paper, we propose a novel two-stage framework DiffusionGAN3D, which boosts text-guided 3D domain adaption and avatar generation by combining the 3D GANs and diffusion priors. Specifically, we integrate the pre-trained 3D generative models (e.g., EG3D) with the text-to-image diffusion models. The former, in our framework, set a strong foundation for text-to-avatar, enabling stable and high-quality 3D avatar generation. In return, the latter provides informative direction for 3D GANs to evolve, which facilitates the text-guided domain adaption of 3D GANs in an efficient way. Moreover, we introduce a progressive texture refinement stage, which significantly enhances the texture quality of the generation results. Extensive experiments demonstrate that the proposed framework achieves excellent results in both domain adaption and text-to-avatar tasks, outperforming existing methods in terms of generation quality and efficiency.

### 12 Apr 2024 v3

# DiffusionGAN3D: Boosting Text-guided 3D Generation and Domain Adaptation by Combining 3D GANs and Diffusion Priors

@@ Abstract
Text-guided domain adaptation and generation of 3D-aware portraits find many applications in various fields. However, due to the lack of training data and the challenges in handling the high variety of geometry and appearance, the existing methods for these tasks suffer from issues like inflexibility, instability, and low fidelity. In this paper, we propose a novel framework DiffusionGAN3D, which boosts text-guided 3D domain adaptation and generation by combining 3D GANs and diffusion priors. Specifically, we integrate the pre-trained 3D generative models (e.g., EG3D) and text-to-image diffusion models. The former provides a strong foundation for stable and high-quality avatar generation from text. And the diffusion models in turn offer powerful priors and guide the 3D generator finetuning with informative direction to achieve flexible and efficient text-guided domain adaptation. To enhance the diversity in domain adaptation and the generation capability in text-to-avatar, we introduce the relative distance loss and case-specific learnable triplane respectively. Besides, we design a progressive texture refinement module to improve the texture quality for both tasks above. Extensive experiments demonstrate that the proposed framework achieves excellent results in both domain adaptation and text-to-avatar tasks, outperforming existing methods in terms of generation quality and efficiency. The project homepage is at https://younglbw.github.io/DiffusionGAN3D-homepage/.
@@ 1 Introduction
Some works [56, 2, 51, 53, 25, 26, 47] perform transfer learning on a pre-trained 3D GAN generator to achieve 3D stylization, which relies on a large number of stylized images and strictly aligned camera poses for training. [2, 47] leverage existing 2D-GAN trained on a specific domain to synthesize training data and implement finetuning with adversarial loss. In contrast, [51, 25, 26] utilize text-to-image diffusion models to generate training datasets in the target domain. This enables more flexible style transferring but also brings problems like pose bias, tedious data processing, and heavy computation costs. Unlike these adversarial finetuning based methods, StyleGAN-Fusion [48] adopts SDS [37] loss as guidance of text-guided adaptation of 2D and 3D generators, which gives a simple yet effective way to fulfill domain adaptation. However, it also suffers from limited diversity and suboptimal text-image correspondence.
@@
Based on the observations above, we propose a novel two-stage framework DiffusionGAN3D to boost the performance of 3D domain adaptation and text-to-avatar tasks by combining 3D generative models and diffusion priors, as shown in Fig. 2. For the text-guided 3D Domain Adaptation task, we first leverage diffusion models and adopt SDS loss to finetune a pre-trained EG3D-based model [7, 4, 9] with random noise input and camera views. The relative distance loss is introduced to deal with the loss of diversity caused by the SDS technique. Additionally, we design a diffusion-guided reconstruction loss to adapt the framework to local editing scenarios. Then, we extend the framework to Text-to-Avatar task by finetuning 3D GANs with a fixed latent code that is obtained guided by CLIP [38] model. During optimization, a case-specific learnable triplane is introduced to strengthen the generation capability of the network. To sum up, in our framework, the diffusion models offer powerful text-image priors, which guide the domain adaptation of the 3D generator with informative direction in a flexible and efficient way. In turn, 3D GANs provide a strong foundation for text-to-avatar, enabling stable and high-quality avatar generation. Last but not least, taking advantage of the powerful 2D synthesis capability of diffusion models, we propose a Progressive Texture Refinement module as the second stage for these two tasks above, which significantly enhances the texture quality. Extensive experiments demonstrate that our method exhibits excellent performance in terms of generation quality and stability on 3D domain adaptation and text-to-avtar tasks, as shown in Fig. 1.
@@
(A) We achieve text-guided 3D domain adaptation in high quality and diversity by combining 3D GANs and diffusion priors with the assistance of the relative distance loss.
@@ 2 Related Work
Domain Adaptation of 3D GANs. The advancements in 3D generative models [7, 35, 4, 9, 6, 11, 14, 13, 29, 45] have enabled geometry-aware and pose-controlled image generation. Especially, EG3D [7] utilizes triplane as 3D representation and integrates StyleGAN2 [24] generator with neural rendering [33] to achieve high-quality 3D shapes and view-consistency image synthesis, which facilitates the downstream applications such as 3D stylization, GAN inversion [28]. Several works [22, 56, 2, 53] achieve 3D domain adaptation by utilizing stylized 2D generator to synthesize training images or distilling knowledge from it. In contrast, [25, 26, 51] leverage the powerful diffusion models to generate training datasets in the target domain and accomplish text-guided 3D domain adaptation with great performance. Though achieving impressive results, these adversarial learning based methods above suffer from issues such as pose bias, tedious data processing, and heavy computation cost. Recently, non-adversarial finetuining methods [3, 12, 48] also exhibit great promise in text-guided domain adaptation. Especially, StyleGAN-Fusion [48] adopts SDS loss as guidance for the adaptation of 2D generators and 3D generators. It achieves efficient and flexible text-guided domain adaptation but also faces the problems of limited diversity and suboptimal text-image correspondence.
@@ 3 Methods
In this section, we present DiffusionGAN3D, which boosts the performance of 3D domain adaptation and text-to-avatar by combining and taking advantage of 3D GANs and diffusion priors. Fig. 2 illustrates the overview of our framework. After introducing some preliminaries (Sec. 3.1), we first elaborate our designs in diffusion-guided 3D domain adaptation (Sec. 3.2) , where we propose a relative distance loss to resolve the problem of diversity loss caused by SDS. Then we extend this architecture and introduce a case-specific learnable triplane to fulfill 3D-GAN based text-to-avatar (Sec. 3.3). Finally, we design a novel progressive texture refinement stage (Sec. 3.4) to improve the detail and authenticity of the texture generated above.
@@ 3.1 Preliminaries
3.2 Diffusion-Guided 3D Domain Adaptation
Due to the difficulties in obtaining high-quality pose-aware data and model training, adversarial learning methods for 3D domain adaptation mostly suffer from the issues of tedious data processing and mode collapse. To address that, we leverage diffusion models and adopt the SDS loss to implement transfer learning on an EG3D-based 3D GAN to achieve efficient 3D domain adaptation, as shown in Fig. 2.
Given a style code $\bm{w}$ generated from noise $\bm{z}\sim N(0,1)$ through the fixed mapping network, we can obtain the triplane $\bm{T}$ and the image $\bm{x}$ rendered in a view controlled by the input camera parameters $\bm{c}$ using the triplane generator and decoder in sequence. Then SDS loss (Sec. 3.1) is applied on $\bm{x}$ to finetune the network. Different from DreamFusion which optimizes a NeRF network to implement single object generation, we shift the 3D generator with random noise and camera pose to achieve domain adaptation guided by text $\bm{y}$. During optimization, all parameters of the framework are frozen except the triplane generator. We find that the gradient provided by SDS loss is unstable and can be harmful to some other well-trained modules such as the super-resolution module. Besides, freezing the mapping network ensures that the latent code $\bm{w}$ lies in the same domain during training, which is a crucial feature that can be utilized in the diversity preserving of the 3D generator.
Relative Distance Loss. The SDS loss provides diffusion priors and achieves text-guided domain adaptation of 3D GAN in an efficient way. However, it also brings the problem of diversity loss as illustrated in [48]. To deal with that, [48] proposes the directional regularizer to regularize the generator optimization process, which improves the diversity to a certain extent. However, it also limits the domain shifting, facing a trade-off between diversity and the degree of style transfer. To address this, we propose a relative distance loss. As shown in Fig. 3, considering two style codes $\bm{w_{i}}$ and $\bm{w_{j}}$ which are mapping from two different noise $\bm{z_{i}}$ and $\bm{z_{j}}$, we project them into the original triplane domain ($\bm{T^{\prime}_{i}}$, $\bm{T^{\prime}_{j}}$) and the finetuned one ($\bm{T_{i}}$, $\bm{T_{j}}$) using a frozen triplane generator $G_{frozen}$ and the finetuned triplane generator $G_{train}$, respectively. Note that, since the mapping network is frozen during training in our framework, $\bm{T_{i}}$ and $\bm{T^{\prime}_{i}}$ (same for $\bm{T_{j}}$ and $\bm{T^{\prime}_{j}}$) share the same latent code and ought to be close in context. Thus, we model the relative distance of these two samples in triplane space and formulate the relative distance loss $L_{dis}$ as:
@@ 3.2 Diffusion-Guided 3D Domain Adaptation
Diffusion-guided Reconstruction Loss. Despite the combination of SDS loss and the proposed relative distance loss is adequate for most domain adaptation tasks, it still fails to handle the local editing scenarios. A naive solution is to perform reconstruction loss between the rendered image and the one from the frozen network. However, it will also inhibit translation of the target region. Accordingly, we propose a diffusion-guided reconstruction loss especially for local editing, which aims to preserve non-target regions while performing 3D editing on the target region. We found that the gradient of SDS loss has a certain correlation with the target area, especially when the noise level $t$ is large, as shown in Fig. 4. To this end, we design a diffusion-guided reconstruction loss $L_{diff}$ that can be presented as:
@@
To sum up, we can form the loss functions for normal domain adaptation and local editing scenario as $L_{adaptation}=L_{sds}+\lambda_{1}L_{dis}$ and $L_{editing}=L_{sds}+\lambda_{2}L_{diff}$, respectively, where $\lambda_{1}$ and $\lambda_{2}$ are the weighting coefficients.
@@ 4.1 Implementation Details
Our framework is built on an EG3D-based model in the first stage. Specifically, we implement 3D domain adaptation on PanoHead, EG3D-FFHQ, and EG3D-AFHQ for head, face, and cat, respectively. For text-to-avatar tasks, PanoHead and AG3D are adopted as the base generators for head and body generation. We employ StableDiffusion v2.1 as our pre-trained text-to-image model. In the texture refinement stage, StableDiffusion v1.5 coupled with ControlNets are utilized to implement image-to-image and inpainting. More details about the parameters and training setting are specified in supplementary materials.
@@ 4.2 Qualitative Comparison
Figure 6: The qualitative comparisons on 3D domain adaptation (applied on EG3D-FFHQ [7]).
For 3D Domain adaptation, we evaluate our model with two powerful baselines: StyleGAN-NADA* [12] and StyleGAN-Fusion [48] for text-guided domain adaptation of 3D GANs, where * indicates the extension of the method to 3D models. For a fair comparison, we use the same prompts as guidance for all the methods. Besides, the visualization results of different methods are sampled from the same random noise. As shown in Fig. 6, the naive extension of StyleGAN-NADA* for EG3D exhibits poor results in terms of diversity and image quality. StyleGAN-Fusion achieves decent 3D domain adaptation and exhibits a certain diversity. However, the proposed regularizer of StyleGAN-Fusion also hinders itself from large-gap domain transfer, resulting in a trade-off between the degree of stylization and diversity. As Fig. 6 shows that the generated faces of StyleGAN-Fusion lack diversity and details, and the hair and clothes suffer from inadequate stylization. In contrast, our method exhibits superior performance in diversity, image quality, and text-image correspondence.
Figure 7: Visual comparisons on text-to-avatar task. The first two rows are the results of ‘head’ and the rest are the results of ‘body’.
@@ 4.3 Quantitative Comparison
We quantitatively evaluate the above baselines and our method on 3D domain adaptation through FID [15] comparison and user study. Specifically, all methods are employed to conduct domain adaptation on EG3D-face and EG3D-cat with both four text prompts, respectively. For each text prompt, we utilize the text-to-image diffusion model to generate 2000 images with different random seeds as the ground truth for FID calculation. In the user study, 12 volunteers were invited to rate each finetuned model from 1 to 5 based on three dimensions: text-image correspondence, image quality, and diversity. As shown in Table  1, the proposed methods achieve lower FID scores than other baselines, which indicates superior image fidelity. Meanwhile, the user study demonstrates that our method outperforms the other two methods, especially in terms of image quality and diversity.
@@
Table 1: Quantitative comparison on 3D domain adaptation task.
@@ 4.4 Ablation Study
On relative distance loss. As shown in Fig. 9, if adopting SDS loss alone for domain adaptation, the generator will tend to collapse to a fixed output pattern, losing its original diversity. In contrast, the proposed relative distance loss effectively preserves the diversity of the generator without sacrificing the stylization degree.
@@
Figure 10: Ablation study of the diffusion guided reconstruction loss. The ToRGB module in EG3D is trained together with $G_{train}$. The input text is “a close-up of a woman with green hair”.
@@ 5 Conclusion
In this paper, we propose a novel two-stage framework DiffusionGAN3D, which boosts text-guided 3D domain adaptation and avatar generation by combining the 3D GANs and diffusion priors. Specifically, we integrate the pre-trained 3D generative models (e.g., EG3D) with the text-to-image diffusion models. The former, in our framework, set a strong foundation for text-to-avatar, enabling stable and high-quality 3D avatar generation. In return, the latter provides informative direction for 3D GANs to evolve, which facilitates the text-guided domain adaptation of 3D GANs in an efficient way. Moreover, we introduce a progressive texture refinement stage, which significantly enhances the texture quality of the generation results. Extensive experiments demonstrate that the proposed framework achieves excellent results in both domain adaptation and text-to-avatar tasks, outperforming existing methods in terms of generation quality and efficiency.